OpenStudy (anonymous):

what two numbers that the square root of 7 is between

OpenStudy (anonymous):

\[4<7<9\] \[\sqrt{4}<\sqrt{7}<\sqrt{9}\] \[2<\sqrt{7}<3\]

OpenStudy (anonymous):

\[\sqrt{7} \approx 2,6457513110645905905016157536393\] \[2< \sqrt{7} <3\] \[2.6< \sqrt{7} <2.7\] \[2.64< \sqrt{7} <2.65\] etc...

OpenStudy (anonymous):

thanx :)

OpenStudy (anonymous):

what about 11 ?

OpenStudy (anonymous):

\[9<11<16\] \[\sqrt{9}<\sqrt{11}<\sqrt{16}\] \[3<\sqrt{11}<4\]